# CompoDiff: Versatile Composed Image Retrieval With Latent Diffusion

### v3

@@ 1 Introduction
We aim to achieve a generalizable CIR model with diverse and versatile conditions by using latent diffusion. We treat the CIR task as a conditional image editing task on the latent space, i.e., $z_{i}=\texttt{Edit}(z_{i_{R}}|z_{c},\ldots)$. Our diffusion-based CIR model, named CompoDiff, can easily deal with versatile and complex conditions, benefiting from the flexibility of the latent diffusion model (Rombach et al., 2022) and the classifier-free guidance (Ho & Salimans, 2022). We train a latent diffusion model that translates the embedding of the reference image ($z_{i_{R}}$) into the embedding of the target image ($z_{i}$) guided by the embedding of the given text condition ($z_{c}$). As shown in Fig. 1, CompoDiff can handle various conditions, which is not possible with the standard CIR scenario with the limited text condition $x_{c_{T}}$. Although our method has an advantage over the existing fusion-based CIR methods in terms of versatility, CompoDiff also needs to be trained with triplet datasets where the scale of the existing CIR datasets is extremely small.
@@
To show the generalizability of the models, we evaluate the models on the “zero-shot” (ZS) CIR scenario using four CIR benchmarks: FashionIQ (Wu et al., 2021), CIRR (Liu et al., 2021), CIRCO (Baldrati et al., 2023), and GeneCIS (Vaze et al., 2023); i.e., we report the retrieval results by the models trained on our SynthTriplets18M and a large-scale image-text paired dataset without access to the target triplet datasets. In all experiments, CompoDiff achieves the best zero-shot performances with significant gaps (See Table 3). Moreover, we observe that the fusion-based approaches solely trained on SynthTriplets18M (e.g., Combiner (Baldrati et al., 2022)) show comparable or outperforming zero-shot CIR performances compared to the previous SOTA zero-shot CIR methods, e.g., Pic2Word (Saito et al., 2023) and SEARLE (Baldrati et al., 2023). Furthermore, we qualitatively observe that the retrieval results of CompoDiff are semantically better than previous zero-shot CIR methods, such as Pic2Word, on a large-scale image database, e.g., LAION-2B.
@@ 2 Related Works
Figure 2: Comparisons of CIR methods. (a) Fusion-based methods, such as ARTEMIS (Delmas et al., 2022) and Combiner (Baldrati et al., 2022) make a fused feature from image feature $z_{i_{R}}$ and text feature $z_{c}$. (b) Inversion-based methods, such as Pic2Word (Saito et al., 2023) and SEARLE (Baldrati et al., 2023), project $z_{i_{R}}$ into the text space, then perform text-to-image retrieval. (c) Our approach applies a diffusion process to $z_{i_{R}}$ with classifier-free guidance by additional text or mask conditions. (b) and (c) use frozen encoders, and (a) usually tunes the encoders.
@@
To solve this problem, recent studies aim to solve CIR tasks in a zero-shot manner. Inversion-based zero-shot CIR methods, such as Pic2Word (Saito et al., 2023) and SEARLE (Baldrati et al., 2023), tackle the problem through text-to-image retrieval tasks where the input image is projected into the condition text – Fig. 2 (b). Note that our zero-shot CIR scenario is slightly different from theirs; our zero-shot CIR denotes that the models are not trained on the target triplet datasets, but trained on our synthetic dataset, SynthTriplets18M, and image-text paired dataset, e.g., LAION-2B (Schuhmann et al., 2022a). Meanwhile, Saito et al. (2023) and Baldrati et al. (2023) use the term zero-shot when the CIR models are trained without a triplet dataset.
@@
[…]
@@ 3.3 Inference
where $\varnothing$ denotes null embeddings, i.e., the empty text (“”) CLIP textual embedding for the text null embedding and an all-zero vector for the image null embedding. One of the advantages of Eq. 4 is the ability to handle various conditions at the same time. When using negative text, we simply replace $\varnothing_{i_{T}}$ with the CLIP text embeddings $c_{T^{-}}$ for the negative text.
Figure 5: Inference condition control. Example retrieval results by varying $w_{I}$ and $w_{T}$ in Eq. 4.
@@ 4 SynthTriplets18M: Massive High-Quality Synthesized Dataset
Table 1: Dataset statistics. $\langle x_{t_{R}},x_{c},x_{t}\rangle$ denotes the triplet of captions, i.e., {original caption, modification instruction, and modified caption}, and $\langle x_{i_{R}},x_{c},x_{i}\rangle$ denotes the CIR triplet of {original image, modification instruction, and modified image}.
Compared to manual dataset collections (Wu et al., 2021; Liu et al., 2021), our approach can easily generate more diverse triplets even if a triplet rarely occurs in reality (See the examples in Fig. 6). Compared to the synthetic dataset of IP2P, our generation process is more scalable due to the keyword-based diverse caption generation process: Our caption triplets are synthesized based on keywords, SynthTriplets18M covers more diverse keywords than IP2P (47k vs. 586k as shown in Table 1). As a result, SynthTriplets18M contains more massive triplets (1M vs. 18M), and CIR models trained on SynthTriplets18M achieve better CIR performances even in the same dataset scale (1M).
@@ 4.5 Dataset Statistics
We also compare SynthTriplets18M, FashionIQ, and CIRR in terms of the token statistics of instructions (i.e., $x_{c}$). Fig. 8 shows that the instruction statistics vary across different datasets. We presume that this is why the zero-shot CIR is still difficult to outperform the task-specific supervised CIR methods.
@@ 5.2 Experiment settings
We also compare a naive editing-based retrieval method: editing an image using IP2P with an instruction and retrieving images using the CLIP image encoder. Note that Pic2Word and SEARLE are trained on the image-only datasets, therefore, it is impossible to train them on SynthTriplets18M. Therefore, we only train the previous fusion-based methods, ARTEMIS and Combiner on SynthTriplets18M for comparison.
@@
[…]
@@
${}^{\dagger}$ ViT-L 24.70 43.70 23.90 53.76 8.72 9.51 10.65 11.16
${}^{\dagger}$ ViT-L 27.51 47.90 24.87 53.80 10.18 11.03 12.72 -
${}^{\dagger}$ ViT-L 25.56 46.23 24.24 53.76 11.68 12.73 14.33 12.31
@@ 5.6 Inference time
One possible drawback of CompoDiff is a relatively slow inference compared to encoder-only methods. However, we confirm that CompoDiff is practically useful with high throughput (120 $\sim$ 230ms) for a single image (Table 4). The table reports throuputs of the comparison methods on ViT-L/14 backbone (Pic2Word, SEARLE, CompoDiff) or RN50 backbone (ARTEMIS, Combiner). One of the advantages of CompoDiff is that we can control the trade-off between retrieval performance and inference time. Note that it is impossible for the other methods. If we need a faster inference time, even with a worse retrieval performance, we can reduce the number of diffusion steps as shown in Table 4. Even with only 5 iterations, our model can produce competitive results. If we use 100 steps, we have a slightly better performance (42.65 vs. 42.33), but a much slower inference time (2.02 sec vs. 0.12 sec). In the experiments, we set the denoising step size to 10.
Step 1 5 10 25 50 75 100 Pic2Word SEARLE ARTEMIS Combiner
Avg(R@10, R@50) 21.52 42.17 42.33 41.42 42.45 42.61 42.65 34.20 35.90 40.61 41.84
Time (sec) 0.02 0.12 0.23 0.56 1.08 1.62 2.02 0.02 0.02 0.005 0.006
Table 4: Performances vs. inference time by varying the number of denoising steps. Numbers are measured on the FashionIQ validation split. CompoDiff (ViT-L) in Table 3 is equivalent to 10 steps, but using 5 steps is a practical altenative. The inference time was measured on a single A100 GPU with a batch size of 1.
@@ 5.7 Analysis
Table 5: Comparisons of design choices for handling textual embeddings on cross-modal retrieval benchmarks. Throughput was measured on 1 A100 GPU with a batch size of 32. For all metrics, higher is better.
CompoDiff does not take textual embeddings $z_{c_{T}}$ as concatenated input tokens of the denoising Transformer but as a condition of cross-attention for the efficient inference. We compare the impact of different design choices for handling textual embeddings. First, we evaluate the Dall-e2 “Prior” model (Ramesh et al., 2022) which converts CLIP textual embeddings into CLIP visual embeddings (we use a public community model  [^4] https://huggingface.co/kakaobrain/karlo-v1-alpha because the official model is not yet publically available). Second, we test the “Prior-like” model by using the denoising Transformer, but taking text guidance as input tokens instead of cross-attention. We also test two more CompoDiff models from our two-stage training strategy.
@@
As observed by Balaji et al. (2022), using a text-oriented model in addition to the CLIP textual encoder results in improved performance of image-text tasks. Motivated by this observation, we also use both the CLIP textual encoder and the language-oriented encoder for extracting the text features of Eq. 3. Table 7 shows the choice of the text encoder affects a lot to the performance on ViT-L/14 CLIP backbone. When the CLIP textual encoder and the T5-XL were used together, the results improved significantly. We suspect that this is because the strong T5 encoder can help the CLIP text encoder to better understand given captions. Interestingly, we observe that using T5 alone degrades the performance even compared to using the CLIP textual encoder alone. We presume this is because T5-XL is specified for long text sequences (e.g., larger than 100 tokens) and text-only data. Meanwhile, SynthTriplets18M has a short average length (See Fig. 8 and 8), which is not specialized by T5. Also, our dataset is based on captions, paired with an image; we also need to consider image information to understand the given caption, but we cannot handle image information alone with T5.
@@ 5.8 Qualitative examples
Finally, it is worth noting that CompoDiff generates a feature belonging to the CLIP visual latent space. It means unCLIP (Ramesh et al., 2022), which decodes a CLIP image feature to an image, can be applied to our composed features. We compare the top-1 retrieval results from LAION and the generated images in Fig. 12 and Appendix C. We use the community version ViT-L unCLIP decoderFootnote 4, by replacing the original Prior module to CompoDiff. As shown in the figures, CompoDiff can manipulate the given input reflecting the given conditions. We believe that incorporating unCLIP into the real-world search engine could potentially improve the user experience by generating images when the desired search results are not available.

### v4

@@ 1 Introduction
We aim to achieve a generalizable CIR model with diverse and versatile conditions by using latent diffusion. We treat the CIR task as a conditional image editing task on the latent space, i.e., $z_{i}=\texttt{Edit}(z_{i_{R}}|z_{c},\ldots)$. Our diffusion-based CIR model, named CompoDiff, can easily deal with versatile and complex conditions, benefiting from the flexibility of the latent diffusion model (Rombach et al., 2022) and the classifier-free guidance (Ho & Salimans, 2022). Note that the purpose of image editing tasks and CIR tasks are different; an image editing model aims to find any image satisfying the changes due to the instruction, while a CIR model aims to find a generalized concept of the given composed queries. For example, assume a picture of a dog on a grass field and an instruction “change the dog to a cat”. Here, even if we assume a perfect image generator can generate an image of a cat on a grass field, it could not achieve a good retrieval performance because it cannot generate all possible cat images in the database. In other words, the purpose of image editing is “precision”, while the purpose of image retrieval is “recall”; it makes the difference between image editing and retrieval. In our experiments, we support this claim by showing directly using an editing model performs much worse than other CIR methods. Therefore, we need a specialized model for CIR rather than directly using the image editing model. In this paper, we tackle the problem by using a latent diffusion model instead of an image-level diffusion model and introducing various retrieval-specific conditions, such as mask conditions. We train a latent diffusion model that translates the embedding of the reference image ($z_{i_{R}}$) into the embedding of the target image ($z_{i}$) guided by the embedding of the given text condition ($z_{c}$). As shown in Fig. 1, CompoDiff can handle various conditions, which is not possible with the standard CIR scenario with the limited text condition $x_{c_{T}}$. Although our method has an advantage over the existing fusion-based CIR methods in terms of versatility, CompoDiff also needs to be trained with triplet datasets where the scale of the existing CIR datasets is extremely small.
@@
To show the generalizability of the models, we evaluate the models on the “zero-shot” (ZS) CIR scenario using four CIR benchmarks: FashionIQ (Wu et al., 2021), CIRR (Liu et al., 2021), CIRCO (Baldrati et al., 2023), and GeneCIS (Vaze et al., 2023); i.e., we report the retrieval results by the models trained on our SynthTriplets18M and a large-scale image-text paired dataset without access to the target triplet datasets. In all experiments, CompoDiff achieves the best zero-shot performances with significant gaps (See Table 3). Moreover, we observe that the fusion-based approaches solely trained on SynthTriplets18M (e.g., Combiner (Baldrati et al., 2022)) show comparable or outperforming zero-shot CIR performances compared to the previous SOTA zero-shot CIR methods, e.g., Pic2Word (Saito et al., 2023), SEARLE (Baldrati et al., 2023) and LinCIR (Gu et al., 2024). Furthermore, we qualitatively observe that the retrieval results of CompoDiff are semantically better than previous zero-shot CIR methods, such as Pic2Word, on a large-scale image database, e.g., LAION-2B (Schuhmann et al., 2022a).
@@ 2 Related Works
Figure 2: Comparisons of CIR methods. (a) Fusion-based methods (e.g., ARTEMIS (Delmas et al., 2022) and Combiner (Baldrati et al., 2022)) make a fused feature from image feature $z_{i_{R}}$ and text feature $z_{c}$. (b) Inversion-based methods (e.g., Pic2Word (Saito et al., 2023), SEARLE (Baldrati et al., 2023) and LinCIR (Gu et al., 2024)) project $z_{i_{R}}$ into the text space, then perform text-to-image retrieval. (c) We apply a diffusion process to $z_{i_{R}}$ with classifier-free guidance by additional conditions. (b) and (c) use frozen encoders, and (a) usually tunes the encoders.
@@
To solve this problem, recent studies aim to solve CIR tasks in a zero-shot manner. Inversion-based zero-shot CIR methods, such as Pic2Word (Saito et al., 2023), SEARLE (Baldrati et al., 2023) and LinCIR (Gu et al., 2024), tackle the problem through text-to-image retrieval tasks where the input image is projected into the condition text – Fig. 2 (b). Note that our zero-shot CIR scenario is slightly different from theirs; our zero-shot CIR denotes that the models are not trained on the target triplet datasets, but trained on our synthetic dataset, SynthTriplets18M, and image-text paired dataset, e.g., LAION-2B (Schuhmann et al., 2022a). Meanwhile, Saito et al. (2023), Baldrati et al. (2023) and Gu et al. (2024) use the term zero-shot when the CIR models are trained without a triplet dataset.
@@
Concurrent with our work, CoVR (Ventura et al., 2024), VDG (Jang et al., 2024), and MagicLens (Zhang et al., 2024) consider constructing training triplets from the existing image datasets, namely, collecting image pairs and synthesizing the modification texts using LLMs. Although they showed empirically good performances, their approaches need real images where it is often very difficult to generate plausible modification instructions between two real images. Meanwhile, our synthetic dataset is more controllable than their datasets. Namely, we can exactly guide the visual difference between two images with the given instructions.
@@
A conventional data collection process for $\langle x_{i_{R}},x_{c},x_{i}\rangle$ is two-staged: collecting candidate reference-target image pairs and manually annotating the modification sentences by human annotators. For example, FashionIQ (Wu et al., 2021) collects the candidate pairs from the same item category (e.g., shirt, dress, and top) and manually annotates the relative captions by crowd workers. CIRR (Liu et al., 2021) gathers the candidate pairs from real-life images from NLVR2 (Suhr et al., 2018). The triplet collection process inevitably becomes expensive, making it difficult to scale CIR datasets. We mitigate this problem by generating a massive synthetic dataset instead of relying on human annotators.
@@ 3.3 Inference
where $\varnothing$ denotes null embeddings, i.e., the empty text (“”) CLIP textual embedding for the text null embedding and an all-zero vector for the image null embedding. $w_{I}$ and $w_{T}$ are weights for controlling the effect of image query or text query, respectively. We can control the degree of the visual similarity ($w_{I}$) or the text instruction ($w_{T}$) by simply adjusting the weight values without any training. One of the advantages of Eq. 4 is the ability to handle various conditions at the same time. When using negative text, we simply replace $\varnothing_{i_{T}}$ with the CLIP text embeddings $c_{T^{-}}$ for the negative text. These advantages are not available by the previous CIR methods without additional training.
Figure 5: Inference condition control by varying $w_{I}$, $w_{T}$ in Eq. 4.
@@ 4 SynthTriplets18M: Massive High-Quality Synthesized Dataset
Table 1: Dataset statistics. $\langle x_{t_{R}},x_{c},x_{t}\rangle$ denotes the triplet of captions, i.e., {original caption, instruction, and modified caption}, and $\langle x_{i_{R}},x_{c},x_{i}\rangle$ denotes the CIR triplet of {original image, instruction, and modified image}.
Compared to manual dataset collections (Wu et al., 2021; Liu et al., 2021), our approach can easily generate more diverse triplets even if a triplet rarely occurs in reality (See the examples in Fig. 6). For example, FashionIQ only contains text instructions for fashion items (e.g., “is blue and has stripes”), but SynthTriplets18M contains more generic and various instructions. Furthermore, manually annotated instructions can be inconsistent and noisy due to the inherent noisiness of crowdsourcing. For example, CIRR has instructions with no information (e.g., “same environment different species”), or instructions which ignore the original information (e.g., “the target photo is of a lighter brown dog walking in white gravel along a wire and wooden fence”). On the other hand, SynthTriplets18M instructions only have modification information. Hence, a model trained on our CIR triplets can learn meaningful and generalizable representations, showing great ZS-CIR performances. Compared to the synthetic dataset of IP2P, our generation process is more scalable due to the keyword-based diverse caption generation process: Our caption triplets are synthesized based on keywords, SynthTriplets18M covers more diverse keywords than IP2P (47k vs. 586k as shown in Table 1). As a result, SynthTriplets18M contains more massive triplets (1M vs. 18M), and CIR models trained on SynthTriplets18M achieve better CIR performances even in the same dataset scale (1M).
@@ 4.5 Dataset Statistics
We also compare SynthTriplets18M, FashionIQ, CIRR, CIRCO, and GeneCIS in the token statistics of instructions (i.e., $x_{c}$). Fig. 8 shows that the instruction statistics vary across different datasets. We presume that this is why the ZS-CIR still has difficulty outperforming the task-specific supervised CIR methods.
@@ 5.2 Experiment settings
We also compare a naive editing-based retrieval method: editing an image using IP2P with an instruction and retrieving images using the CLIP image encoder. Note that Pic2Word and SEARLE are trained on the image-only datasets; therefore, it is impossible to train them on SynthTriplets18M. Similarly, Combiner and ARTEMIS only take triplets for training; hence, they cannot be trained with image-caption datasets, such as Conceptual Captions (Sharma et al., 2018) or LAION (Schuhmann et al., 2022a). Due to this reason, we only train the previous fusion-based methods, ARTEMIS and Combiner on SynthTriplets18M for comparison. Note that Combiner needs a pre-trained CLIP model trained on a large-scale image-caption dataset. In this point of view, we can argue that Combiner and CompoDiff are comparable in terms of the dataset scale; both methods use a large-scale image-caption dataset and synthetic triplets in our experiment.
@@
CIRR (Liu et al., 2021) contains more generic images than FashionIQ. CIRR uses the images from NLVR2 (Suhr et al., 2018) with more complex and long descriptions. CIRR has 36k open-domain triplets divided into the train, validation, and test sets in 8:1:1 split. The examples of $x_{c}$ are “remove all but one dog and add a woman hugging it”, “It’s a full pizza unlike the other with the addition of spice bottles”, or “Paint the counter behind the brown table white”. As the example captions show, one of the main drawbacks of CIRR is that the text queries are not realistic compared to the real user scenario; we may need more shorter and direct instructions. As the text instruction distribution of CIRR is distinct from others (Fig. 8), we observe that sometimes CIRR results are a not reliable measure of open-world CIR tasks (e.g., retrieval on LAION). Baldrati et al. (2023) also observe another main drawback of CIRR. CIRR contains a lot of false negatives (FNs) due to the nature of the image-text dataset (Chun et al., 2021; 2022; Chun, 2023)
@@
Pic2Word ViT-L 24.70 43.70 23.90 53.76 8.72 9.51 10.65 11.16
SEARLE-OTI ViT-L 27.51 47.90 24.87 53.80 10.18 11.03 12.72 -
SEARLE ViT-L 25.56 46.23 24.24 53.76 11.68 12.73 14.33 12.31
@@ 5.6 Inference time
One possible drawback of CompoDiff is a relatively slow inference compared to encoder-only methods. However, we confirm that CompoDiff is practically useful with high throughput (120 $\sim$ 230ms) for a single image (Table 4). The table reports throuputs of the comparison methods on ViT-L/14 backbone (Pic2Word, SEARLE, CompoDiff) or RN50 backbone (ARTEMIS, Combiner). One of the advantages of CompoDiff is that we can control the trade-off between retrieval performance and inference time. Note that it is impossible for the other methods. If we need a faster inference time, even with a worse retrieval performance, we can reduce the number of diffusion steps as shown in Table 4. Even with only 4 or 5 iterations, our model can produce competitive results. If we use 100 steps, we have a slightly better performance (42.65 vs. 42.17), but a much slower inference time (2.02 sec vs. 0.12 sec). In the experiments, we set the step size to 10.
Step 1 2 3 4 5 10 50 100 Pic2Word SEARLE ARTEMIS Combiner
Avg(R@10, R@50) 21.52 29.22 37.63 41.08 42.17 42.33 42.45 42.65 34.20 35.90 40.61 41.84
Time (sec) 0.02 0.05 0.07 0.09 0.12 0.23 1.08 2.02 0.02 0.02 0.005 0.006
GPU Memory (bs 1) 2.60 GB 0.97 GB 1.32 GB 0.26 GB 0.54 GB
GPU Memory (bs 256) 5.80 GB 2.99 GB 3.24 GB 2.82 GB 1.95 GB
Table 4: Performances vs. inference time by varying the number of denoising steps. Numbers are measured on the FashionIQ validation split. CompoDiff (ViT-L) in Table 3 is equivalent to 10 steps, but using 5 steps is a practical alternative. The inference time was measured on a single A100 GPU with a batch size of 1. GPU memories are the pick-allocated memory measured on batch sizes of 1 and 64, respectively. Note that ARTEMIS cannot support efficient batch operation for inference because its forward path needs triplet information, not an image-instruction pair; here, we report ARTEMIS information with training triplets.
@@
We also remark that recently, there have been huge advances in boosting diffusion model inference speed, such as the Consistency Model (Song et al., 2023) or distillation (Meng et al., 2023). Using more efficient diffusion model variants for the CompoDiff model will also greatly improve the inference speed. However, these approaches will need a completely different training framework (Song et al., 2023) or a complicated two-staged distillation process (Meng et al., 2023). While we believe that applying these techniques to CompoDiff will bring a huge benefit, we leave further improvements for future work.
@@ 5.7 Analysis
Table 5: Comparisons of design choices for handling textual embeddings on cross-modal retrieval benchmarks. Throughput was measured on 1 A100 GPU with a batch size of 32. For all metrics, higher is better. Here, we use the community version of the prior model by Lee et al. (2022) because the official prior model (Ramesh et al., 2022) is not public.
CompoDiff does not take textual embeddings $z_{c_{T}}$ as concatenated input tokens of the denoising Transformer but as a condition of cross-attention for the efficient inference. We compare the impact of different design choices for handling textual embeddings. First, we evaluate the Dall-e2 “Prior” model (Ramesh et al., 2022) which converts CLIP textual embeddings into CLIP visual embeddings (we use a public community model (Lee et al., 2022) because the official model is not yet publically available). Second, we test the “Prior-like” model by using the denoising Transformer, but taking text guidance as input tokens instead of cross-attention. We also test two more CompoDiff models from our two-stage training strategy.
@@
As observed by Balaji et al. (2022) and Byun et al. (2024), the power of the CLIP textual encoder affects a lot to the performance of image-text tasks. Motivated by this observation, we also use both the CLIP textual encoder and the language-oriented encoder for extracting the text features of Eq. 3. Table 7 shows the choice of the text encoder affects a lot to the performance on ViT-L/14 CLIP backbone. When the CLIP textual encoder and the T5-XL were used together, the results improved significantly. We suspect that this is because the strong T5 encoder can help the CLIP text encoder to better understand given captions. Interestingly, we observe that using T5 alone degrades the performance even compared to using the CLIP textual encoder alone. We presume this is because T5-XL is specified for long text sequences (e.g., larger than 100 tokens) and text-only data. Meanwhile, SynthTriplets18M has a short average length (See Fig. 8 and 8), which is not specialized by T5. Also, our dataset is based on captions, paired with an image; we also need to consider image information to understand the given caption, but we cannot handle image information alone with T5.
@@ 5.8 Qualitative examples
Finally, it is worth noting that CompoDiff generates a feature belonging to the CLIP visual latent space. Namely, unCLIP (Ramesh et al., 2022), which decodes a CLIP image feature to an image, can be applied to our composed features. We compare the top-1 retrieval results from LAION and the generated images in Fig. 12 and Appendix C. We use the Karlo unCLIP decoder (Lee et al., 2022), by replacing the original Prior module to CompoDiff. As shown in the figures, CompoDiff can manipulate the given input reflecting the given conditions. We believe that incorporating unCLIP into the real-world search engine could potentially improve the user experience by generating images when the desired search results are not available.